The last submission was incorrect, this is it. Can somene show me how to use the intercepts to graph the equation
x + 3y = 6

Answer by rfer(16322)

y intercept
0+3y=6
y=6/3
y=2
(0,2)
x intercept
x+3*0=6
x+0=6
x=6
(6,0)